"On June 30, 2017, Macrosoft Company acquired a 10-acre tract of land. On the tract was a warehouse that Macrosoft intended to use as a distribution center. At the time of purchase, the land had an assessed tax value of $6,300,000, and the building had an assessed tax value of $11,200,000. Macrosoft paid $15,000,000 for the land and building. After the purchase, the company paid $1,000,000 to have various modifications made to the building.Required: 1. At what amount should Macrosoft record the land and building? 2. For financial reporting purposes, why might Macrosoft managers prefer to assign a large"
Answer #1

Answer 1 :- At what amount should Macrosoft record the land and building ?

Land :-

Calculation of land portion from total FMV of land and building portion

FMV of Land / FMV of Land & Building = 6,300,000 / (6,300,000+11,200,000)

= 6,300,000 / 17,500,000

= 36%

Cost of Land = 15,000,000 * 36% = $5,400,000

Building :-

Calculation of building portion from total FMV of land and building portion

FMV of building / FMV of Land & Building = 11,200,000 / (6,300,000+11,200,000)

= 11,200,000 / 17,500,000

= 64%

Cost of building = (15,000,000 * 64%) + Cost of modification of building

= 9,600,000 + 100,000

= $10,600,000

Answer 2:-.

Land is non depreciable assets so no depreciation is recorded on land. If the higher amount is allocated to the land means will be lower depreciation in future year and this will result higher net income of such years.